# EDITOR: Effective and Interpretable Prompt Inversion for Text-to-Image Diffusion Models

### v1

@@ 4 Experiment Setup
Datasets and Models. We evaluate our method on three widely-used datasets: MS COCO [17], which contains diverse natural images with detailed annotations; LAION [18], a large-scale dataset of image-text pairs; and Flickr [19], which consists of diverse images, including everyday life scenes, natural landscapes, animals, urban settings, and human activities. For each dataset, we randomly sample 100 images for evaluation. We employ Stable Diffusion v1.5 [3] as our target generative model for prompt inversion experiments.
@@
  • Image Similarity: We utilize CLIP [38] and LPIPS [39] scores to measure the visual similarity between generated and target images, assessing the reconstruction quality.
  • Textual Alignment: BERTScore [40] is used to evaluate semantic consistency between inverted and ground-truth prompts, computing Precision, Recall, and F1 metrics.
  • Prompt Interpretability: We calculate Perplexity Scores (PPL) to assess the linguistic coherence of inverted prompts, where lower scores indicate more natural and comprehensible text.
Baseline Methods. We compare our approach with state-of-the-art prompt inversion methods: PEZ [12], which inverts prompts with image-text similarity matching, and PH2P [11], which employs a delayed projection scheme to optimize prompts.
@@ 5 Experiment Results
Table 4: Images generated with inverted prompts. For a given target image(left), inverted prompts from PEZ [12], PH2P [11] and our method are used to prompt Stable Diffusion (v1.5) [3], DALL-E 3 [41] and Ideogram 2.0 (right).
@@
Image Similarity. We evaluate the effectiveness of the inverted prompts generated by EDITOR for image generation. For a given target image, we first perform inversion to recover the corresponding prompt and assess the relevance of the images generated by the diffusion model using the recovered prompt. We compare EDITOR with PEZ [12] and PH2P [11] under different datasets. The results are shown in subsection 3.5. For CLIP similarity, EDITOR achieves the highest scores. In detail, the average CLIP similarity score under MSCOCO, LAION and Flickr are 0.794, 0.826 and 0.776, respectively, outperforming PEZ and PH2P. Furthermore, EDITOR achieves a value of 0.414 under MSCOCO, 0.401 under LAION and 0.424 under Flickr in terms of the LPIPS similarity, consistently outperforming the baselines. The results demonstrate that EDITOR’s prompts achieve high reconstruction quality in terms of similarity to the target images.
Textual Alignment. We evaluate the semantic similarity between the inverted prompts and the ground-truth prompts using BERTScore [40]. As shown in subsection 3.5, EDITOR achieves consistently higher precision, recall, and F1 scores compared to the baseline methods across all datasets. Notably, under MS COCO, LAION and Flickr, EDITOR achieves a precision of 0.900, 0.841, and 0.892, surpassing the baselines by a value of 0.099, 0.047 and 0.094. When considering recall, EDITOR also excels across datasets. In detail, EDITOR can achieve a recall score of 0.920, 0.824 and 0.915 under all three datasets, respectively, outperforming the baselines. These results illustrate the effectiveness of our approach in achieving semantically meaningful text recovery that aligns closely with the ground-truth prompts.
Prompt Interpretability. We further assess the interpretability of the recovered prompts by computing their perplexity scores (PPL). Lower perplexity indicates that the inverted prompts are more fluent, natural, and interpretable. Results in subsection 3.5 show that our approach achieves significantly lower perplexity scores than PEZ and PH2P. Specifically, EDITOR achieves a PPL score of 80.659 under MS COCO, 187.299 under LAION, and 90.320 under Flickr, which far outperform the scores of PEZ (8,837.797 under MS COCO, 8,008.401 under LAION, 9,425.353 under Flickr) and PH2P (11,078.994 under MS COCO, 9,621.671 under LAION, 11,804.100 under Flickr). On MS COCO and Flickr, EDITOR shows 110% and 104% improvement compared to the baselines, while on LAION, it demonstrates 42% better performance. The vast gap in performance highlights the exceptional ability of EDITOR to recover prompts that are not only semantically aligned with the ground-truth but also significantly more fluent and interpretable.
Comparison with Image Captioning Models. To demonstrate the superior performance of EDITOR compared to state-of-the-art image captioning models or vision large language models including LLaVa-1.5-7B, Janus-Pro-7B, BLIP-Large, BLIP2-OPT-2.7B, and GIT, we evaluated their differences in image generation quality. As shown in Table 5, EDITOR achieves the highest CLIP Similarity score (0.776) and the lowest LPIPS Similarity score (0.424) when compared to other image captioning models like LLaVa-1.5-7B [14], Janus-Pro-7B [15], BLIP-large [42], BLIP2-OPT-2.7B [13], and GIT-large [43]. This indicates that the prompts generated by EDITOR have a stronger connection to the visual content of the images and are more consistent with human-labeled captions.
Table 5: Comparison with advanced image captioning models or vLLMs.
Metric EDITOR LLaVA-1.5-7B Janus-Pro-7B BLIP-Large BLIP2-OPT-2.7B GIT
@@
An important characteristic of EDITOR is its ability to optimize captions generated by an image captioning model in conjunction with a target diffusion model. This leads to improved text-image alignment. Figure 4 shows the enhancements achieved using EDITOR across datasets. The results clearly demonstrate that EDITOR significantly enhances the similarity score between generated images and the original images, increasing the CLIP similarity by at least 5%. Figure 5 shows the improvements achieved using EDITOR across different image captioning models. The results indicate that EDITOR can significantly improve the quality of generated images across various captioning models. What’s more, we can fully believe that if image understanding models or captioning models have nice performance, EDITOR can achieve better results based on them.
Ablation Studies. To show the importance of initialization in EDITOR, we compared random initialization and image captioning model initialization on a Flickr subset. As shown in section 6, random prompts hinder convergence, often leading to poor performance and local optima traps. Our initialization is crucial for guiding the algorithm to accurate, optimal solutions. For fairness, we also initialized PEZ and PH2P. As shown in section 6, initialization only slightly improved PEZ’s textual Alignment and prompt interpretability, with almost no impact on other aspects for both methods. This is because PEZ and PH2P are discrete optimization algorithms; even if they start with interpretable prompts, the optimization process may gradually replace them with high-perplexity token combinations. This also demonstrates that our method has greater potential than existing baselines, particularly in its compatibility with initialization. For additional ablation studies, refer to Appendix C.
Qualitative Results. We show example inverted prompts towards four target images and corresponding generations with three models: Stable Diffusion [3], DALL-E 3 [41], and Ideogram 2.0 in  Table 4. EDITOR produces more accurate, readable prompts that generate images closely aligned with the target images, demonstrating strong generalizability across models. In contrast, prompts from PEZ and PH2P tend to be less coherent, often resulting in images that do not align well with the intended concepts, highlighting the superior precision and versatility of our approach.
6 Application of the generated prompts
Figure 6: Application of cross-concept image synthesis.
Figure 7: Application of removal or replacement of objects.
Removal or Replacement of Concepts. Our prompt inversion method enables removal or replacement of image concepts. Unlike PEZ [12] and PH2P [11], which produce hard-to-interpret prompts, our method generates clear, human-readable prompts, simplifying token identification for concept editing. This improves flexibility in image manipulation. For example, Figure 7 shows removing "trees" from the prompt eliminates them, while replacing "trees" with "fence" substitutes a fence, demonstrating effective prompt-based editing. For more applications including evolutionary multi-concept generation and unsupervised segmentation, refer to Appendix B.
@@ 7 Conclusion
In this paper, we introduced a prompt inversion approach for text-to-image diffusion models that generates high-quality, semantically aligned and interpretable prompts. Our method, validated on datasets like MS COCO, LAION, and Flickr, outperforms existing techniques in image similarity, textual alignment, and efficiency. Additionally, the generated prompts are versatile and applicable to tasks such as cross-concept image synthesis and object removal or replacement. These results demonstrate that prompt inversion is a promising technique for improving the performance and versatility of text-to-image generation models.
Appendix A Details about the E2T Model
@@
For the embedding-to-text (E2T) model, we use T5-base [44] as the backbone, and the training was performed on the MSMARCO corpus [45] with a batch size of 32, a learning rate of $1\times 10^{-3}$, a maximum token length of 32, an epoch number of 60 for the zero-step model and an epoch number of 35 for the correction model. We test the E2T model using 2,500 prompts selected from the MS COCO dataset [17], with evaluation metrics based on BertScore [40], including Precision (P), Recall (R), and F1-score. The results was shown in Table 8. Our E2T model achieves over 0.968 in Precision, Recall, and F1-score, demonstrating that our model can accurately reverse the embedding back into the original prompt with high precision.
Appendix B Applications of Prompt Inversion
@@
Removal or Replacement of Concepts. Our prompt inversion method can also be applied to the removal or replacement of concepts within an image. Unlike PEZ [12] and PH2P [11], which often generate prompts that are difficult to interpret, making it challenging to identify the corresponding tokens for the objects to be replaced, our method generates prompts that are fully human-readable and understandable. This clarity allows for easy identification of the specific tokens associated with the concepts, facilitating the seamless replacement or removal of those concepts within the image. This advantage enhances the flexibility and usability of our approach in image manipulation tasks. As shown in Figure 9, removing the word "egg" from the prompt eliminates them from the image, while replacing "egg" with "broccoli" substitutes the egg with a broccoli. This demonstrates the flexibility of our method in editing images via prompt modification.
@@
Unsupervised Segmentation. EDITOR can be applied to unsupervised semantic segmentation by leveraging cross-attention maps to create segmentation masks [46, 47]. EDITOR generates prompts capturing key concepts directly from a diffusion model, without external data. For a target image, we use EDITOR to produce the inverted prompt, and as shown in 10, the resulting cross-attention maps align tokens with relevant image regions, enabling accurate segmentation.
@@
Evolutionary Multi-concept Generation. Evolutionary multi-concept image generation is a process that enables users to create complex images by iteratively combining concepts from multiple sample images using prompt inversion. Consider a example in 11, we used EDITOR to get two prompts from two images: "a giraffe burger with bacon and cheese and french fries" and "there is a bowl of spinach and some tomatoes in it." We then combined these two prompts to create a new image, which resulted in a new prompt through EDITOR: "Burger with fries on a plate with salad and tomatoes on it.".
@@
Appendix C Ablation studies
Impact of Initial Prompt Length. Since our prompt inversion method starts with an initial prompt, we explore the impact of the initial prompt length on the inversion performance. We vary the initial prompt length from 12 to 20 tokens, and the results are shown in Appendix C. The results show that an initialization length of 16 tokens achieves the best performance. Shorter lengths (e.g., 12 tokens) tend to lack sufficient semantic information, making it harder to accurately reconstruct prompts. In contrast, longer prompts (e.g., 20 tokens) provide richer semantic information, leading to improved textual alignment and prompt interpretability. For instance, EDITOR achieves a prompt precision of 0.884 with 20 tokens, surpassing the precision of 0.852 achieves with 12 tokens. However, longer prompts introduce excessive complexity, making the optimization process more challenging. Compared with 20 tokens, the result of 16 tokens achieves a higher CLIP Similarity of 0.789 and better performance in textual alignment and prompt interpretability. At the same time, this length also offers a lower optimization complexity. As a result, an appropriate initialization length such as 16 tokens could be selected to balance semantic richness and optimization efficiency.

### v2

@@ 4 Experiment Setup
Datasets. We evaluate our method on three widely-used datasets: MS COCO (Lin et al., 2014), which contains diverse natural images with detailed annotations; LAION (Schuhmann et al., 2022), a large-scale dataset of image-text pairs; and Flickr (Young et al., 2014), which consists of diverse images, including everyday life scenes, natural landscapes, animals, urban settings, and human activities. For each dataset, we randomly sample 100 images for evaluation.
Models. We employ Stable Diffusion v1.5 (Rombach et al., 2022) as our target text-to-image generative model for prompt inversion experiments. In addition, we compare our method against different image captioning models and vision language models, including LLaVA-1.5-7B (Liu et al., 2024), Janus-Pro-7B (Chen et al., 2025), BLIP-Large (Li et al., 2022b), BLIP2-OPT-2.7B (Li et al., 2023), and GIT-Large (Wang et al., 2022), and further conduct experiments on advanced multi-encoder architectures including SDXL-Turbo (Sauer et al., 2023) and Stable Diffusion 3.5 Medium (Esser et al., 2024). More detailed descriptions of these models can be found in the Appendix A.
@@
  • Image Similarity: We utilize CLIP (Radford et al., 2021a) and LPIPS (Zhang et al., 2018) scores to measure the visual similarity between generated and target images, assessing the reconstruction quality.
  • Textual Alignment: BERTScore (Zhang et al., 2020) is used to evaluate semantic consistency between inverted and ground-truth prompts, computing Precision, Recall, and F1 metrics.
  • Prompt Interpretability: We calculate Perplexity(PPL) scores to assess the linguistic coherence of inverted prompts, where lower scores indicate more comprehensible text.
Baseline Methods. We compare our approach with state-of-the-art prompt inversion methods: PEZ (Wen et al., 2024), which inverts prompts with image-text similarity matching, PH2P (Mahajan et al., 2024), which employs a delayed projection scheme to optimize prompts for text-to-image diffusion models and VGD (Kim et al., 2025), which uses a large language model plus CLIP-based guidance to generate coherent, human-readable prompts.
@@ 5 Experiment Results
Table 4: Images generated with inverted prompts. For a given target image (left), inverted prompts from PEZ Wen et al. (2024), PH2P Mahajan et al. (2024) and our method are used to prompt Stable Diffusion v1.5 Rombach et al. (2022), DALL-E 3 Betker et al. (2023) and Ideogram 2.0 (right).
@@
Table 5: Comparison with advanced image captioning models or vision language models.
Metric EDITOR LLaVA-1.5-7B Janus-Pro-7B BLIP-Large BLIP2-OPT-2.7B GIT-Large
@@
Table 6: Evaluation on advanced multi-encoder models.
CLIP Score $\uparrow$ LPIPS Score $\downarrow$ Prompt Precision $\uparrow$ Prompt Recall $\uparrow$ Prompt F1 $\uparrow$ Prompt PPL $\downarrow$
Stable Diffusion v1.5 0.829 0.426 0.842 0.825 0.834 154.362
SDXL Turbo 0.792 0.428 0.824 0.818 0.821 105.114
Stable Diffusion 3.5 Medium 0.785 0.445 0.814 0.817 0.815 115.132
@@
[…]
Image Similarity. We evaluate the effectiveness of the inverted prompts generated by EDITOR for image generation. For a given target image, we first perform inversion to recover the corresponding prompt and assess the relevance of the images generated by the diffusion model using the recovered prompt. We compare EDITOR with PEZ, PH2P and VGD under different datasets. The results are shown in  Table 3. For CLIP score, EDITOR achieves the highest scores. In detail, the average CLIP similarity score under MSCOCO, LAION and Flickr are 0.794, 0.826 and 0.776, respectively, outperforming PEZ ,PH2P and VGD. Furthermore, EDITOR achieves a value of 0.414 under MSCOCO, 0.401 under LAION and 0.424 under Flickr in terms of the LPIPS score, consistently outperforming the baselines. The results demonstrate that EDITOR’s prompts achieve high reconstruction quality in terms of similarity to the target images.
Textual Alignment. We evaluate the semantic similarity between the inverted prompts and the ground-truth prompts using BERTScore (Zhang et al., 2020). As shown in  Table 3, EDITOR achieves consistently higher precision, recall, and F1 scores compared to the baseline methods across all datasets. Notably, under MS COCO, LAION and Flickr, EDITOR achieves a precision of 0.900, 0.841, and 0.892, surpassing the baselines by a value of 0.074, 0.041 and 0.066. When considering recall, EDITOR also excels across datasets. In detail, EDITOR can achieve a recall score of 0.920, 0.824 and 0.915 under all three datasets, respectively, outperforming the baselines. These results illustrate the effectiveness of our approach in achieving semantically meaningful text recovery that aligns closely with the ground-truth prompts.
Prompt Interpretability. We further assess the interpretability of the recovered prompts by computing their perplexity scores (PPL). Lower perplexity indicates that the inverted prompts are more fluent, natural, and interpretable. Results in  Table 3 demonstrates that EDITOR shows 110% and 104% improvement compared to PEZ and PH2P on MS COCO and Flickr. Moreover, even when compared with VGD, which leverages large language models to generate fluent prompts with relatively low perplexity (665.418 under MS COCO, 700.754 under LAION, and 536.712 under Flickr), our method still achieves much lower scores. The vast gap in performance highlights the exceptional ability of EDITOR to recover prompts that are not only semantically aligned with the ground-truth but also significantly more fluent and interpretable.
Comparison with Image Captioning Models. To demonstrate the superior performance of EDITOR compared to state-of-the-art image captioning models or vision language models, we evaluated their differences in image generation quality. As shown in Table 5, EDITOR achieves the highest CLIP Similarity score (0.776) and the lowest LPIPS Similarity score (0.424) when compared to other image captioning models. This indicates that the prompts generated by EDITOR have a stronger connection to the visual content of the images and are more consistent with human-labeled captions.
Performance on Multi-Encoder Diffusion Models. In addition to single-encoder diffusion models, we also evaluate EDITOR on advanced multi-encoder models, which integrate multiple text encoders to enhance text-image alignment. As shown in Table 6, EDITOR consistently achieves strong results across different model architectures. For instance, EDITOR obtains a CLIP score of 0.792 on SDXL-Turbo and 0.785 on Stable Diffusion 3.5 Medium, both close to the performance of the single-encoder Stable Diffusion v1.5 . Moreover, the generated prompts maintain low perplexity, indicating that the outputs remain interpretable and linguistically coherent even under more complex architectures. These results suggest that EDITOR is not limited to a specific model design and can generalize well across both single-encoder and multi-encoder diffusion models, highlighting its robustness and broad applicability.
Figure 5: Application of cross-concept image synthesis.
Figure 6: Application of removal or replacement of objects.
Ablation Studies. To show the importance of initialization in EDITOR, we compared random initialization and image captioning model initialization on a Flickr subset. As shown in Table 5, random prompts hinder convergence, often leading to poor performance and local optima traps. Our initialization is crucial for guiding the algorithm to accurate, optimal solutions. For fairness, we also initialized PEZ and PH2P. As shown in Table 5, initialization only slightly improved PEZ’s textual Alignment and prompt interpretability, with almost no impact on other aspects for both methods. This is because PEZ and PH2P are discrete optimization algorithms; even if they start with interpretable prompts, the optimization process may gradually replace them with high-perplexity token combinations. This also demonstrates that our method has greater potential than existing baselines, particularly in its compatibility with initialization.
Qualitative Results. We show example inverted prompts towards four target images and corresponding generations with three models: Stable Diffusion  (Rombach et al., 2022), DALL-E 3 (Betker et al., 2023), and Ideogram 2.0 in  Table 4. EDITOR produces more accurate, readable prompts that generate images closely aligned with the target images, demonstrating strong generalizability across models. In contrast, prompts from PEZ and PH2P tend to be less coherent, often resulting in images that do not align well with the intended concepts, highlighting the superior precision and versatility of our approach.
6 Application of the generated prompts
Removal or Replacement of Concepts. Our prompt inversion method enables removal or replacement of image concepts. Unlike PEZ (Wen et al., 2024) and PH2P (Mahajan et al., 2024), which produce hard-to-interpret prompts, our method generates clear, human-readable prompts, simplifying token identification for concept editing. This improves flexibility in image manipulation. For example, Figure 6 shows removing "trees" from the prompt eliminates them, while replacing "trees" with "fence" substitutes a fence, demonstrating effective prompt-based editing. For more applications including evolutionary multi-concept generation and unsupervised segmentation, refer to Appendix D.
@@ 7 Conclusion
In this paper, we introduced a prompt inversion approach for text-to-image diffusion models that generates high-quality, semantically aligned and interpretable prompts. Our method, validated on datasets like MS COCO, LAION, and Flickr, outperforms existing techniques in image similarity, textual alignment, and prompt interpretability. Additionally, the generated prompts are versatile and applicable to tasks such as cross-concept image synthesis and object removal or replacement. These results demonstrate that prompt inversion is a promising technique for improving the performance and versatility of text-to-image generation models.
Ethics Statement
As the economic and creative value of prompts in text-to-image generation models grows, there are rising ethical concerns regarding prompt inversion techniques. Our method is capable of recovering high-quality prompts from generated images, which raises potential risks of prompt stealing. Such misuse could infringe upon the intellectual property of prompt engineers and undermine the commercial ecosystem of the prompt marketplace (Shen et al., 2024; Naseh et al., 2024). We acknowledge this concern and emphasize that our work is intended to advance the understanding of this emerging issue and to provide insights into potential safeguards for protecting prompt intellectual property, rather than to facilitate malicious exploitation.
In addition, our method can be applied to data attribution and model provenance. While these applications have positive implications for accountability and transparency, they also highlight broader ethical considerations about privacy, fairness, and legal compliance when tracing data origins. To mitigate risks, we encourage responsible use of EDITOR within controlled research and verification settings. We are committed to ensuring that this research contributes constructively to the dialogue on intellectual property protection, responsible AI, and ethical deployment of generative models.
Appendix A More Details of the Used Models
In this section, we provide more details about the model used in the experiments.
We conduct experiments on three representative text-to-image generative models with different architectural designs, including Stable Diffusion v1.5(Rombach et al., 2022), SDXL Turbo(Sauer et al., 2023), and Stable Diffusion 3.5 Medium(Esser et al., 2024).
  • Stable Diffusion v1.5111https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5. This model is initialized with the weights of the Stable-Diffusion-v1.2222https://huggingface.co/CompVis/stable-diffusion-v1-2 checkpoint and subsequently fine-tuned on 595k steps at resolution 512x512 on "laion-aesthetics v2 5+" and 10% dropping of the text-conditioning to improve classifier-free guidance sampling. The text encoder of this model is CLIP-ViT/L.
  • SDXL Turbo [^3] https://huggingface.co/stabilityai/sdxl-turbo. This model is a distilled version of SDXL 1.0444https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, trained for real-time synthesis. SDXL-Turbo is based on a novel training method called Adversarial Diffusion Distillation (ADD), which allows sampling large-scale foundational image diffusion models in 1 to 4 steps at high image quality. This approach uses score distillation to leverage large-scale off-the-shelf image diffusion models as a teacher signal and combines this with an adversarial loss to ensure high image fidelity even in the low-step regime of one or two sampling steps. The text encoders of this model are OpenCLIP-ViT/G and CLIP-ViT/L.
  • Stable Diffusion 3.5 Medium [^5] https://huggingface.co/stabilityai/stable-diffusion-3.5-medium. This model is a Multimodal Diffusion Transformer with improvements (MMDiT-X) text-to-image model that features improved performance in image quality, typography, complex prompt understanding, and resource-efficiency. The text encoders of this model are OpenCLIP-ViT/G, CLIP-ViT/L and T5-xxl.
In addition to text-to-image diffusion models, we also compare our method with several representative image captioning models and vision-language models that are widely used for generating natural language descriptions from images, including LLaVA-1.5-7B (Liu et al., 2024), Janus-Pro-7B (Chen et al., 2025), BLIP-large (Li et al., 2022b), BLIP2-OPT-2.7B (Li et al., 2023) and GIT-large (Wang et al., 2022).
  • LLaVA-1.5-7B [^6] https://huggingface.co/llava-hf/llava-1.5-7b-hf. This model is an open-source chatbot trained by fine-tuning LlamA/Vicuna on GPT-generated multimodal instruction-following data. It is an auto-regressive language model, based on the transformer architecture. In other words, it is an multi-modal version of LLMs fine-tuned for chat/instructions.
  • Janus-Pro-7B [^7] https://huggingface.co/deepseek-ai/Janus-Pro-7B. This model is a novel autoregressive framework that unifies multimodal understanding and generation. It addresses the limitations of previous approaches by decoupling visual encoding into separate pathways, while still utilizing a single, unified transformer architecture for processing. The decoupling not only alleviates the conflict between the visual encoder’s roles in understanding and generation, but also enhances the framework’s flexibility. Janus-Pro surpasses previous unified model and matches or exceeds the performance of task-specific models. The simplicity, high flexibility, and effectiveness of Janus-Pro make it a strong candidate for next-generation unified multimodal models.
  • BLIP-Large [^8] https://huggingface.co/Salesforce/blip-image-captioning-large. This model is a vision-language pretraining (VLP) framework designed for both understanding and generation tasks. Most existing pretrained models are only good at one or the other. It uses a captioner to generate captions and a filter to remove the noisy captions. This increases training data quality and more effectively uses the messy web data.
  • BLIP2-OPT-2.7B [^9] https://huggingface.co/Salesforce/blip2-opt-2.7b. This model is initialized with weights of the image encoder and large language model from pre-trained checkpoints and keep them frozen while training the Querying Transformer, which is a BERT-like Transformer encoder that maps a set of "query tokens" to query embeddings, which bridge the gap between the embedding space of the image encoder and the large language model. The goal for the model is simply to predict the next text token, giving the query embeddings and the previous text.
  • GIT-Large [^10] https://huggingface.co/microsoft/git-large. This model is a transformer decoder conditioned on both CLIP image tokens and text tokens. The model is trained using "teacher forcing" on a lot of (image, text) pairs. The goal for the model is simply to predict the next text token, giving the image tokens and previous text tokens.
Appendix B Improving Image Captioning with EDITOR
An important characteristic of EDITOR is its ability to optimize captions generated by an image captioning model in conjunction with a target diffusion model. This leads to improved text-image alignment. Figure 7 shows the enhancements achieved using EDITOR across datasets. The results clearly demonstrate that EDITOR significantly enhances the similarity score between generated images and the original images, increasing the CLIP similarity by at least 5%. Figure 8 shows the improvements achieved using EDITOR across different image captioning models. The results indicate that EDITOR can significantly improve the quality of generated images across various captioning models. What’s more, we can fully believe that if image captioning models or vision-language models have nice performance, EDITOR can achieve better results based on them.
Appendix C Details about the E2T Model
@@
For the embedding-to-text (E2T) model, we use T5-base [^11] https://huggingface.co/google-t5/t5-base (Raffel et al., 2020a) as the backbone, and the training was performed on the MSMARCO corpus (Bajaj et al., 2016) with a batch size of 32, a learning rate of $1\times 10^{-3}$, a maximum token length of 32, an epoch number of 60 for the zero-step model and an epoch number of 35 for the correction model. The E2T model learns a general mapping from the diffusion model’s embedding space to natural language, and does not introduce any additional information about the prompt inversion task. This training process is completely decoupled from the main prompt inversion task and its datasets.
We test the E2T model using 2,500 prompts selected from the MS COCO dataset (Lin et al., 2014), with evaluation metrics based on BertScore (Zhang et al., 2020), including Precision (P), Recall (R), and F1-score. The results was shown in Table 9. Our E2T model achieves over 0.968 in Precision, Recall, and F1-score, demonstrating that our model can accurately reverse the embedding back into the original prompt with high precision.
Appendix D Applications of Prompt Inversion
@@
Removal or Replacement of Concepts. Our prompt inversion method can also be applied to the removal or replacement of concepts within an image. Unlike PEZ Wen et al. (2024) and PH2P Mahajan et al. (2024), which often generate prompts that are difficult to interpret, making it challenging to identify the corresponding tokens for the objects to be replaced, our method generates prompts that are fully human-readable and understandable. This clarity allows for easy identification of the specific tokens associated with the concepts, facilitating the seamless replacement or removal of those concepts within the image. This advantage enhances the flexibility and usability of our approach in image manipulation tasks. As shown in Figure 10, removing the word "egg" from the prompt eliminates them from the image, while replacing "egg" with "broccoli" substitutes the egg with a broccoli. This demonstrates the flexibility of our method in editing images via prompt modification.
@@
Unsupervised Segmentation. EDITOR can be applied to unsupervised semantic segmentation by leveraging cross-attention maps to create segmentation masks Wu et al. (2023); Chefer et al. (2023). EDITOR generates prompts capturing key concepts directly from a diffusion model, without external data. For a target image, we use EDITOR to produce the inverted prompt, and as shown in Figure 11, the resulting cross-attention maps align tokens with relevant image regions, enabling accurate segmentation.
@@
Evolutionary Multi-concept Generation. Evolutionary multi-concept image generation is a process that enables users to create complex images by iteratively combining concepts from multiple sample images using prompt inversion. Consider a example in Figure 12, we used EDITOR to get two prompts from two images: "a giraffe burger with bacon and cheese and french fries" and "there is a bowl of spinach and some tomatoes in it." We then combined these two prompts to create a new image, which resulted in a new prompt through EDITOR: "Burger with fries on a plate with salad and tomatoes on it.".
@@
Appendix E Ablation studies
Impact of Initial Prompt Length. Since our prompt inversion method starts with an initial prompt, we explore the impact of the initial prompt length on the inversion performance. We vary the initial prompt length from 12 to 20 tokens, and the results are shown in Table E. The results show that an initialization length of 16 tokens achieves the best performance. Shorter lengths (e.g., 12 tokens) tend to lack sufficient semantic information, making it harder to accurately reconstruct prompts. In contrast, longer prompts (e.g., 20 tokens) provide richer semantic information, leading to improved textual alignment and prompt interpretability. For instance, EDITOR achieves a prompt precision of 0.884 with 20 tokens, surpassing the precision of 0.852 achieves with 12 tokens. However, longer prompts introduce excessive complexity, making the optimization process more challenging. Compared with 20 tokens, the result of 16 tokens achieves a higher CLIP Similarity of 0.789 and better performance in textual alignment and prompt interpretability. At the same time, this length also offers a lower optimization complexity. As a result, an appropriate initialization length such as 16 tokens could be selected to balance semantic richness and optimization efficiency.